# Sequential Regression Learning with Randomized Algorithms

### v1

@@ Abstract.
This paper presents “randomized SINDy”, a sequential machine learning algorithm designed for dynamic data that has a time-dependent structure. It employs a probabilistic approach, with its PAC learning property rigorously proven through the mathematical theory of functional analysis. The algorithm dynamically predicts using a learned probability distribution of predictors, updating weights via gradient descent and a proximal algorithm to maintain a valid probability density. Inspired by SINDy ([SINDy]), it incorporates feature augmentation and Tikhonov regularization. For multivariate normal weights, the proximal step is omitted to focus on parameter estimation. The algorithm’s effectiveness is demonstrated through experimental results in regression and binary classification using real-world data.
@@ 1. Introduction
[…]
2. Sequential Machine Learning
In this section, we will introduce the sequential machine learning model. Consider $(\chi,\beta(\chi))$ a Borel space corresponding to the domain space, in which $\beta(\chi)$ corresponds to the $\sigma$-Borel algebra of the subsets of $\chi$. Consider $(\mathcal{Y},\beta(\mathcal{Y}))$ a Borel space corresponding to the space of labels. We need to develop a learning rule capable of predicting the future label of the process $y\in\mathcal{Y}$ based on the input data $x\in\chi$. This rule is given by a measurable Borel function $h:\chi\rightarrow\mathcal{Y}$. This function is called a predictor or a hypothesis or a classifier. The predictor will be used to predict new labels based on the input points $(x\in\chi)$. We denote ¯:=$\bar{\mathcal{H}}:=\{h:\chi\rightarrow\mathcal{Y},~{}\mbox{Measurable Borel}\}$ as the class of possible predictors and $\mathcal{H}\subset\bar{\mathcal{H}}$ a subclass of admissible predictors. We assume that $\mathcal{H}$ is also a Borel space equipped with the $\sigma$-algebra $\beta(\mathcal{H})$. We denote ${\bf o}_{n}=\{(x_{1},y_{1}),(x_{2},y_{2}),\cdots,(x_{n},y_{n})\}\in\mathbb{O}^{n}$ as the sequence of labelled points corresponding to the training data. This is the input we have available to conduct the learning process.
In the sequel, we will introduce the dynamics of the sequential machine learning system. Here, we consider that the data is generated by a probability distribution with any dependency structure and that the learning algorithms are randomized. In this case, the learner takes an initial action $A_{1}$ which is a probability over the space of admissible predictors $(\mathcal{H},\beta(\mathcal{H}))$. We denote $\mathcal{D}_{1}\subset\mathcal{P}(\mathcal{H})$ as the class of all the learner’s admissible algorithms in Step 1. Next, we observe the pair $(x_{1},y_{1})$ with joint probability $\nu_{1}$ defined over the space of observations $(\mathbb{O},\beta(\mathbb{O}))$. In the proper version of sequential machine learning, the learner’s loss is given by a function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow[0,1]$ Borel measurable. At the end of Stage 1, the learner receives the following loss
$J_{1}({\bf o}_{1},A_{1}):=\int_{\mathcal{H}}\ell(y_{1},h(x_{1}))A_{1}(dh),\quad{\bf o}_{1}\in\mathbb{O}~{}~{}\text{e}~{}~{}A_{1}\in\mathcal{D}_{1}.$
In the second stage of the game, the learner’s action is a Borel measurable transition probability $A_{2}:\mathbb{O}\rightarrow\mathcal{P}(\mathcal{H})$. We denote $\mathcal{D}_{2}$ as the class of all the learner’s admissible algorithms in Step 2. Next, we observe the pair $(x_{2},y_{2})$ with Borel measurable transition probability $\nu_{2}:\mathbb{O}\rightarrow\mathcal{P}(\mathbb{O})$. At the end of Step 2, the learner receives the following loss
$J_{2}({\bf o}_{2},A_{2})=\int_{\mathcal{H}}\ell(y_{2},h(x_{2}))A_{2}(dh\mid o_{1}),\quad{\bf o}_{2}\in\mathbb{O}^{2}~{}~{}\text{e}~{}~{}A_{2}\in\mathcal{D}_{2}.$
In stage $t$ of the game, the learner’s action is a Borel measurable transition probability $A_{t}:\mathbb{O}^{t-1}\rightarrow\mathcal{P}(\mathcal{H})$. We denote $\mathcal{D}_{t}$ as the class of all the learner’s admissible algorithms at stage $t$. Next we observe the pair $(x_{t},y_{t})$ with Borel measurable transition probability $\nu_{t}:\mathbb{O}^{t-1}\rightarrow\mathcal{P}(\mathbb{O})$. At the end of stage $t$, the learner receives the following loss
$J_{t}({\bf o}_{t},A_{t})=\int_{\mathcal{H}}\ell(y_{t},h(x_{t}))A_{t}(dh\mid{\bf o}_{t-1}),\quad{\bf o}_{t}\in\mathbb{O}^{t}~{}~{}\text{e}~{}~{}A_{t}\in\mathcal{D}_{t}.$
We also denote $\mathcal{D}^{t}=\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{t}$ as the set of admissible strategies up to time $t$. The joint probability for nature’s strategy up to stage $t$ is given by:
$\mathbb{P}_{t}(D_{1}\times\cdots\times D_{t})=\int_{D_{1}}\cdots\int_{D_{t}}\nu_{t}(do_{n}\mid{\bf o}_{t-1})\cdots\nu_{1}(do_{1}),\quad D_{i}\in\beta(\mathbb{O}),~{}i=1,\cdots,t.$
The cost function related to the $t$-th step of the sequential machine learning problem is given by the regret function
(2.1) $\begin{split}&R_{t}(\mathbb{P}_{t},{\bf A}_{t}):=\int_{\mathbb{O}^{t}}\left[\sum_{s=1}^{t}J_{s}({\bf o}_{s},A_{s})-\inf_{A\in\mathcal{P}(\mathcal{H})}\sum_{s=1}^{t}\int_{\mathcal{H}}\ell\left(y_{s},h(x_{s})\right)A(dh)\right]\mathbb{P}_{t}(d{\bf o}_{t})\end{split}$
for every randomized algorithm ${\bf A}_{t}=(A_{1},\cdots,A_{t})\in\mathcal{D}^{t}$ and $t\in\mathbb{N}$. Based on the PAC learning property, we will use the following sequential machine learning concept.
Definition 2.1.
A $\mathcal{H}$ class of admissible predictors admits the sequential PAC property if there is a function $m^{PACS}:(0,1)\rightarrow\mathbb{N}$ and a family of learning algorithms ${\bf A}_{t}=(A_{1},\cdots,A_{t})\in\mathcal{D}^{t}$ such that
(2.2) $\sup_{\mathbb{P}_{t}\in\mathbb{P}(\mathbb{O}^{t})}\frac{1}{t}R_{t}(\mathbb{P}_{t},{\bf A}_{t})\leq\delta,\quad t\geq m^{PACS}(\delta),$
for every $\delta\in(0,1)$.
3. Absolutely Continuous Random Algorithm
In order to deduce a PAC learning algorithm, we assume that the learner’s actions are defined by absolutely continuous probabilities with respect to a measure $\delta$ in which the Radon-Nikodym derivatives are square integrable. We denote the space of absolutely continuous probabilities over $(\mathcal{H},\beta(\mathcal{H}))$ and the class of admissible strategies respectively by
$\mathcal{P}_{\delta}(\mathcal{H}):=\left\{\gamma\in\mathcal{P}(\mathcal{H}):\gamma<<\delta~{}~{}\text{e}~{}~{}\frac{\partial\gamma}{\partial\delta}\in L^{2}(\mathcal{H},\delta)\right\}$
and
$\mathcal{D}_{\delta,t}:=\{A_{t}\in\mathcal{D}_{t}:A_{t}(\cdot\mid{\bf o}_{t-1})\in\mathcal{P}_{\delta}(\mathcal{H}),~{}{\bf o}_{t-1}\in\mathbb{O}^{t-1}\}.$
Thus, we can identify strategies as densities with respect to the $\delta$ measure. For every pair of $o=(x,y)\in\mathbb{O}$, we take
$c_{(x,y)}(g):=\int_{\mathcal{H}}\ell(y,h(x))g(h)\delta(dh),$
where $g$ is the density of the algorithm with respect to the measure $\delta$. As a consequence, we consider the problem of minimizing the following functional
(3.1) $\min_{g\in\mathcal{P}_{\delta}(\mathcal{H})}c_{(x,y)}(g)=\min_{g\in L^{2}(\mathcal{H},\delta)}\left\{c_{(x,y)}(g)+\Psi(g)\right\},$
where
$\Psi(g)=\begin{cases}0&;~{}~{}~{}g\in\mathcal{P}_{\delta}(\mathcal{H})\\ \infty&;~{}~{}~{}g\not\in\mathcal{P}_{\delta}(\mathcal{H}).\end{cases}$
As $\mathcal{P}_{\delta}(\mathcal{H})$ is a closed convex subset of $L^{2}(\mathcal{H},\delta)$, we conclude that $\Psi:L^{2}(\mathcal{H},\delta)\rightarrow\mathbb{R}\cup\{\infty\}$ is proper, convex and lower semicontinuous. On the other hand, $c_{(x,y)}:L^{2}(\mathcal{H},\delta)\rightarrow\mathbb{R}$ is a continuously differentiable convex function. In fact, the Gâteaux derivative at point $g$ and direction $\tilde{g}$ is given by
$\nabla c_{(x,y)}\left(g\right)\tilde{g}$ $=$ $\lim_{s\downarrow 0}\frac{c_{(x,y)}(g+s\tilde{g})-c_{(x,y)}(g)}{s}=\int_{\mathcal{H}}\ell(y,h(x))\tilde{g}(h)\delta(dh)$
$=$ $\langle\tilde{g},\ell(y,\cdot(x))\rangle_{L^{2}(\mathcal{H},\delta)}=\langle\tilde{g},\nabla c_{(x,y)}\left(g\right)\rangle_{L^{2}(\mathcal{H},\delta)}.$
Then, $\nabla c_{(x,y)}(g)$ is a continuous linear functional, which can be identified with an element of the Hilbert space $L^{2}(\mathcal{H},\delta)$ (Riesz representation). In this case, we obtain from the section 3 that the functional $\nabla c_{(x,y)}(g)$ is represented by $\ell(y,\cdot(x))$. As the gradient $\nabla c_{(x,y)}$ is constant, we conclude that $\nabla c_{(x,y)}$ is Lipschitz with constant $L>0$. In this case, we have
$c_{(x,y)}(g)=c_{(x,y)}(\theta)=\int_{\mathcal{H}}\ell(y,h(x))g(h,\theta)\delta(dh)=\mathbb{E}_{\theta}\ell(y,h(x))$
Finally, the fact that $c_{(x,y)}+\Psi$ is a coercive function implies that $\operatorname*{arg\,min}\{c_{(x,y)}+\Psi\}$ is nonempty (see, Theorem 2.19 in [Peypouquet]). To make easy notation, we denote $c_{(x,y)}(g)+\Psi(g)$ by $\Theta_{x,y}(g)$ for all $g\in L^{2}(\mathcal{H},\delta)$ and $g^{\star}\in\operatorname*{arg\,min}_{g}\{\Theta\}$. In the sequel, we will derive the forward and backward proximal algorithm (FBA). We take $g_{t-1}\in\mathcal{P}_{\delta}(\mathcal{H})$ and a parameter $\eta_{t-1}>0$. The update of the learner’s action $g_{t}$ is the only point of minimum of the Moreau-Yosida regularization given by
(3.3) $g_{t}:=\operatorname*{arg\,min}\left\{\Theta_{(x,y)}(g)+\frac{1}{2\eta_{t-1}}\parallel g-g_{t-1}\parallel_{L^{2}}^{2}:g\in L^{2}(\mathcal{H},\delta)\right\}.$
According to the Moreau-Rockafellar theorem, we have
(3.4) $0=\nabla\Theta_{(x,y)}(g_{t})+\frac{g_{t}-g_{t-1}}{\eta_{t-1}}=\nabla c_{(x,y)}(g_{t})+\partial\Psi(g_{t})+\frac{g_{t}-g_{t-1}}{\eta_{t-1}},$
where $\partial\Psi$ is a subgradient of $\Psi$ at the point $g_{t}$. Hence, we conclude that
$g_{t}=\left(g_{t-1}-\eta_{t-1}\nabla c_{(x,y)}(g_{t})\right)-\eta_{t-1}\partial\Psi(g_{t})$
or
$\left(I+\eta_{t-1}\partial\Psi\right)(g_{t})=g_{t-1}-\eta_{t-1}\nabla c_{(x,y)}(g_{t})=(I-\eta_{t-1}\nabla c_{(x,y)}(g_{t})(g_{t-1}),$
where $I:L^{2}(\mathcal{H},\delta)\rightarrow L^{2}(\mathcal{H},\delta)$ is the identity operator and also
$g_{t}$ $=$ $\left[\left(I+\eta_{t-1}\partial\Psi\right)^{-1}(I-\eta_{t-1}\nabla c_{(x,y)}(g_{t})\right](g_{t-1})$
$=$ $\operatorname*{prox}_{\eta_{t-1}\Psi}\left[g_{t-1}-\eta_{t-1}\nabla c_{(x,y)}(g_{t})\right]$
where
$\operatorname*{prox}_{\eta\Psi}(v)=\operatorname*{arg\,min}\left\{\Psi(z)+\frac{1}{2\eta}\parallel z-v\parallel^{2}:~{}z\in L^{2}(\mathcal{H},\delta)\right\}.$
The sequence $\{g_{t}\}$ generated by the proximal algorithm is called a proximal sequence. The stationary points of the proximal sequence minimize the objective function $\Theta_{(x,y)}$, because $g_{t}=g_{t-1}$ if, and only if, $\nabla\Theta_{(x,y)}=0$.
Lemma 3.1.
For all $g\in L^{2}(\mathcal{H},\delta)$ and ${\bf o}_{n}\in\mathbb{O}^{n}$, we have
$2\sum_{t=2}^{n+1}\eta_{t-1}\left[c_{(x_{t},y_{t})}(g_{t})-c_{(x_{t},y_{t})}(g)\right]\leq\parallel g_{1}-g^{\star}\parallel_{L^{2}}$
where $\eta_{t}>0$ for all $t=1,2\cdots,n$ and $n\geq 1$.

### v2

# Randomized SINDy

@@ Abstract.
This paper presents “randomized SINDy”, a sequential machine learning algorithm designed for dynamic data that has a time-dependent structure. It employs a probabilistic approach, with its PAC learning property rigorously proven through the mathematical theory of functional analysis. The algorithm dynamically predicts using a learned probability distribution of predictors, updating weights via gradient descent and a proximal algorithm to maintain a valid probability density. Inspired by SINDy ([2]), it incorporates feature augmentation and Tikhonov regularization. For multivariate normal weights, the proximal step is omitted to focus on parameter estimation. The algorithm’s effectiveness is demonstrated through experimental results in regression and binary classification using real-world data.
@@ 1. Introduction
Sequential machine learning represents a family of efficient and scalable algorithms to incrementally build predictive models from a sequence of data ([14]). Traditional machine learning algorithms have a high computational cost for retraining in the presence of new data while sequential algorithms are suitable for applications with a high volume of data sampled dynamically and quickly. Huge databases that grow dynamically over time require scalable machine learning algorithms [9, 10, 11].
Traditional machine learning algorithms assume that the data is independent and identically distributed (iid). However, in many cases involving the mining of text, voice and time series data, the data in step $t$ depends on the data from the previous steps. This brings us to sequential learning algorithms. Recurrent neural networks (RNN) are a class of machine learning algorithms used to treat data correlated in time as time series. However, RNN have a high computational cost, which makes them difficult to apply in the sequential learning scenario [15, 5].
@@
Furthermore, these general predictive models do not provide us with a measure of confidence in the forecasts they make. In various applications, such as demand forecasting, it is important to establish a measure of confidence in the forecasts so that we can develop strategies for stock control. One strategy for developing a measure of confidence in forecasts is to use Bayesian techniques to quantify the uncertainty related to the model and estimates [13, 3]. Another alternative is to use “Conformal Prediction” developed by [16], which uses the assumption of independent observations to construct a set in the space of labels with a given probability [3]. Although the conformal prediction technique is distribution-free, it assumes that the data is independent.
The main objective of this work is to develop a class of algorithms that works incrementally that is suitable for large volumes of data updated online and has a time-dependent structure. In addition, we will propose a measure of confidence for the predictions made by the algorithm. To do this, we develop an extension of the class of SINDy models proposed by [2], incorporating random coefficients that are estimated and updated online. In this case, we leverage online convex optimization techniques to sequentially update the coefficients and to obtain probabilistic coverage guarantees for the prediction estimate.
This paper is organized as follows: in section 2, we introduce the proposed algorithm, randomized SINDy, and demonstrates its PAC learning properties through rigorous analysis. In section 3, we present experimental results, starting with simple simulations to verify the effectiveness of the proposed approach, followed by applications on two real-world datasets. Finally, section 4 concludes with a discussion of the results and future research directions.
2. Randomized SINDy
We consider the problem of sequential learning by the following online convex optimization setting. At the initial step, we observe $x_{1}\in\mathbb{R}^{k}$. Then the learner takes an initial action $A_{1}$ which is a probability distribution over the space of labels $\mathbb{R}^{m}$ as follows
$A_{1}(x_{1},B_{1}):=\mathbb{E}_{b_{1}}\left(1\!\!1_{B_{1}}\Delta(x_{1})\beta_{1}\right),\quad B_{1}\in\beta(\mathbb{R}^{m}),$
where $\Delta(x)\in\mathbb{R}^{m\times p}$, for every $x\in\mathbb{R}^{k}$, is a feature library and $\beta_{1}$ has a $p$-dimensional multivariate distribution with density $g_{b_{1}}$ parametrized by $b_{1}\in\Theta\subset\mathbb{R}^{u}$. We assume that $\Theta$ is a bounded convex and compact set in the Euclidean space. We denote by $D$ an upper bound on the diameter of $\Theta$:
$\parallel b_{1}-b_{2}\parallel\leq D,\quad\forall~b_{1},b_{2}\in\Theta.$
As a consequence, the action $A_{1}$ can be identified by the parameter $b_{1}\in\Theta$. Next, we observe $y_{1}\in\mathbb{R}^{m}$ and the learner receives the following loss
$L(b_{1},(x_{1},y_{1}))=\mathbb{E}_{b_{1}}\left(\ell(y_{1},\Delta(x_{1})\beta_{1})+\lambda\beta_{1}^{T}\beta_{1}\right)$
where $\lambda\geq 0$ is the complexity parameter and $\ell:\mathbb{R}^{m}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$ is an integrable function. At the second step, we observe $x_{2}\in\mathbb{R}^{k}$. Then the learner takes an action $A_{2}$ which is a probability distribution over the space of labels $\mathbb{R}^{m}$ as follows
$A_{2}(x_{2},B_{2}):=\mathbb{E}_{b_{2}}\left(1\!\!1_{B_{2}}\Delta(x_{2})\beta_{2}\right),\quad B_{2}\in\beta(\mathbb{R}^{m}),$
where $\beta_{2}$ has a $p$-dimensional multivariate distribution with density $g_{b_{2}}$ parametrized by $b_{2}\in\Theta$. As a consequence, the action $A_{2}$ can be identified by the parameter $b_{2}\in\Theta$. Next, we observe $y_{2}\in\mathbb{R}^{m}$ and the learner receives the following loss
$L(b_{2},(x_{2},y_{2}))=\mathbb{E}_{b_{2}}\left(\ell(y_{2},\Delta(x_{2})\beta_{2})+\lambda\beta_{2}^{T}\beta_{2}\right).$
Following in this way, at the step $t$, we observe $x_{t}\in\mathbb{R}^{k}$. Then the learner takes an action $A_{t}$ which is a probability distribution over the space of labels $\mathbb{R}^{m}$ as follows
$A_{t}(x_{t},B_{t}):=\mathbb{E}_{b_{t}}\left(1\!\!1_{B_{t}}\Delta(x_{t})\beta_{t}\right),\quad B_{t}\in\beta(\mathbb{R}^{m}),$
where $\beta_{t}$ has a $p$-dimensional multivariate distribution with density $g_{b_{t}}$ parametrized by $b_{t}\in\Theta$. As a consequence, the action $A_{t}$ can be identified by the parameter $b_{t}\in\Theta$. Next, we observe $y_{t}\in\mathbb{R}^{m}$ and the learner receives the following loss
$L(b_{t},(x_{t},y_{t}))=\mathbb{E}_{b_{t}}\left(\ell(y_{t},\Delta(x_{t})\beta_{t})+\lambda\beta_{t}^{T}\beta_{t}\right).$
The goal of a sequential learning algorithm is to find a sequence $\{\hat{b}^{t}:t\geq 1\}$ that achieve the minimum regret along the whole learning process. The regret is defined as
$Regret=\sum_{t=1}^{n}L(\hat{b}^{t}({\bf o}_{t-1}),(x_{t},y_{t}))-\min_{b\in\Theta}\sum_{t=1}^{n}L(b,(x_{t},y_{t})).$
In order to establish that the proximal gradient algorithm minimizes the regret, we suppose that the following hypothesis is valid.
Sequential convexity Hypothesis:
  • For every fixed $(x,y)\in\mathbb{R}^{k}\times\mathbb{R}^{m}$, the function $L(\cdot,(x,y)):\Theta\rightarrow\mathbb{R}$ is given by
    $L(b,(x,y))=\mathbb{E}_{b}\left(\ell(y,\Delta(x)\beta)+\lambda\beta^{T}\beta\right)$
    where $\beta$ has a $p$-dimensional multivariate distribution with density $g_{b}$ parametrized by $b\in\Theta$, is a convex and differentiable function.
  • $\parallel\nabla L(b,(x,y)\parallel\leq G$ with constant $G>0$.
As a consequence, we propose the following sequential proximal gradient algorithm
$\hat{b}^{t}({\bf o}_{t-1})=prox_{\eta_{t-1},\Psi}\left[\hat{b}^{t-1}({\bf o}_{t-2})-\eta_{t-1}\nabla L(\hat{b}^{t-1},(x_{t-1},y_{t-1}))\right]$
where $\hat{b}^{1}\in\Theta$, $t\geq 2$ and
$\operatorname*{prox}_{\eta,\Psi}(v)=\arg\min_{z}\left\{\Psi(z)+\frac{1}{2\eta}\parallel z-v\parallel^{2}\right\},$
with
$\Psi(b)=\begin{cases}0&;~~~b\in\Theta\\ \infty&;~~~b\not\in\Theta.\end{cases}$
Next, we apply usual arguments from the theory of convex learning to prove that the sequential proximal gradient minimizes the regret.
Theorem 2.1.
Under the sequential convexity hypothesis, the regret function satisfies
$\frac{1}{n}\left(\sum_{t=1}^{n}L(\hat{b}^{t}({\bf o}_{t-1}),(x_{t},y_{t}))-\min_{b\in\Theta}\sum_{t=1}^{n}L(b,(x_{t},y_{t}))\right)\leq 3DG\frac{1}{\sqrt{n}},$
with $\eta_{t}=\frac{D}{G\sqrt{t}}$.